# Revisiting Node Affinity Prediction in Temporal Graphs

### v2

@@ 4.1 Node Affinity Prediction on TGB
Table 1: NDCG@10 on TGBN datasets (↑ higher is better). NAVIS is benchmarked against TGNNs that use all available graph messages. Boldface marks the best method.
@@
Table 2: NDCG@10 on TGBN datasets using only previous ground-truth labels (↑ higher is better). This setting is suited for heuristics. Boldface marks the best method. Baselines have no standard deviation because they are pre-defined and deterministic.
@@ 4.2 Generalization to Link Prediction Datasets
To answer (RQ2), we repurpose four temporal link prediction datasets (Wikipedia(Kumar et al., 2019), Flights(Strohmeier et al., 2021), USLegis(Fowler, 2006), and UNVote(Voeten et al., 2009)) for the future node affinity prediction task. We detail how we adjust these datasets in Appendix˜B. TGNNs are known to operate well on these datasets (Yu et al., 2023), and, therefore, should constitute a strong baseline. As shown in Tables 3 and 4, NAViS consistently outperforms both TGNN and heuristic baselines, with gains ranging from +13.9% to +20% over the second performing TGNN, suggesting NAViS generalizes well to many dynamic graph datasets. Similar to the results in Tables˜1 and 2, other TGNNs fall behind simple heuristics even though they were shown to produce great results on these datasets for future link prediction tasks, suggesting an incompatible design choice of TGNNs for future node affinity prediction.
@@ 4.3 Ablation study
To answer (RQ3), we conduct an ablation study to isolate the contribution of each core component of NAViS: (1) the linear state update mechanism (vs. the commonly used GRU), (2) the inclusion of the global virtual vector $\mathbf{g}$}, and (3) the proposed loss (vs. cross-entropy, CE). Full results are provided in Appendix˜E. The ablations show that each component contributes substantially to the overall performance.
@@ 5 Related Work
Heuristics and State Space Models. A key observation in recent literature is that simple heuristics like Moving Average often outperform complex TGNNs on various relevant benchmarks (Huang et al., 2023; Cornell et al., 2025), suggesting the problem is fundamentally sequential. Recent work formally establishes the equivalence between linear SSMs and Moving Average (Eliasof et al., 2025) and explores their use in language modeling (Gu et al., 2022; Gu & Dao, 2024) and dynamic link prediction (Li et al., 2024; Ding et al., 2025). While these works connect SSMs to temporal data, our approach is distinct. We are the first to explicitly leverage the formal equivalence between heuristics and SSMs to design an architecture, NAViS, that is purpose-built for node affinity prediction.
@@ Appendix A Datasets statistics and description
  • tgbn-reddit (Nadiri & Takes, 2022) models Reddit as a temporal bipartite graph linking users and subreddits. Nodes represent both entities, and a timestamped edge records a user’s post within a subreddit. The dataset covers activity from 2005 through 2019. The predictive objective is, for each user, to produce a next-week ranking of subreddits by expected engagement intensity.
@@ Appendix D Implementation Details
We initialized each linear layer in NAViS with values drawn uniformly from the range $\sqrt{d},\sqrt{d}]$, where $d$ is the input dimension of the layer. We adopted standard hyperparameters, consistent with prior work: batch size 200, Adam optimizer, and learning rate $10^{-4}$. Following the TGB protocol, all models were trained for 50 epochs, and the checkpoint with the best performance on the validation set was selected. We set the regularization margin to $\Delta=10^{-3}$. During training, NAViS computes the loss only over the top-20 affinities, ensuring that loss computation requires constant time and memory. We used a single NVIDIA GeForce RTX 3090 GPU and a single AMD Ryzen 9 7900X 12-Core Processor CPU.
Appendix E Ablation of Design Choices
@@
We further ablate on the regularization term of our loss, and performed an additional experiment on the TGB datasets where the regularization term is not included in the loss computation. We report the results in Table˜8.

### v3

@@ 4.1 Node Affinity Prediction on TGB
Table 1: NDCG@10 on TGB datasets (↑ higher is better). NAVIS is benchmarked against TGNNs that use all available graph messages. Boldface marks the best method.
@@
[…]
Table 2: NDCG@10 on TGB datasets using only previous ground-truth labels (↑ higher is better). This setting is suited for heuristics. Boldface marks the best method. Baselines have no standard deviation because they are pre-defined and deterministic.
@@ 4.2 Generalization to Link Prediction Datasets
To answer (RQ2), we repurpose four temporal link prediction datasets (Wikipedia(Kumar et al., 2019), Flights(Strohmeier et al., 2021), USLegis(Fowler, 2006), and UNVote(Voeten et al., 2009)) for the future node affinity prediction task. We detail how we adjust these datasets in Appendix˜B. TGNNs are known to operate well on these datasets (Yu et al., 2023), and, therefore, should constitute a strong baseline. As shown in Tables 3 and 4, NAViS consistently outperforms both TGNN and heuristic baselines, with gains ranging from +13.9% to +20% over the second performing TGNN, suggesting NAViS generalizes well to many dynamic graph datasets. Similar to the results in Tables˜1 and 2, other TGNNs fall behind simple heuristics even though they were shown to produce great results on these datasets for future link prediction tasks, strengthening that the TGNNs design choices mentioned earlier in Sections˜3.1 and 3.3 are incompatible for future node affinity prediction.
@@ 4.3 Ablation study
To answer (RQ3), we conduct an ablation study to isolate the contribution of each core component of NAViS: (1) the linear state update mechanism (vs. the commonly used GRU), (2) the inclusion of the global virtual vector $\mathbf{g}$}, and (3) the proposed loss (vs. cross-entropy, CE). Full results are provided in Appendix˜F. The ablations show that each component contributes substantially to the overall performance.
@@ 5 Related Work
Heuristics and State Space Models. According to the recent literature, simple heuristics, like Moving Average, often outperform complex TGNNs on various relevant benchmarks (Huang et al., 2023; Cornell et al., 2025), suggesting the problem is fundamentally sequential. Recent work formally establishes the equivalence between linear SSMs and Moving Average (Eliasof et al., 2025) and explores their use in language modeling (Gu et al., 2022; Gu and Dao, 2024) and dynamic link prediction (Li et al., 2024; Ding et al., 2025). While these works connect SSMs to temporal data, our approach is distinct. We are the first to explicitly leverage the formal equivalence between heuristics and SSMs to design an architecture, NAViS, that is purpose-built for node affinity prediction.
Temporal link prediction on weighted dynamic graphs. Temporal link prediction (TLP) on weighted dynamic graphs (Qin et al., 2023; Yang et al., 2019; Lei et al., 2019) is the task in which, given a discrete-time weighted dynamic graph (a sequence of snapshots of the graph at specific points in time), the model is required to predict the next snapshot, i.e., the weighted adjacency matrix at a future time. Although this task resembles node affinity prediction, there are several key differences. First, the downstream objectives differ, and accordingly different metrics are used to measure performance in each setting. In TLP on weighted dynamic graphs, the end goal is to construct the entire future weighted adjacency matrix. Hence, metrics such as RMSE, MSE, and MAE are often used. In node affinity prediction, however, the goal is to rank different nodes with respect to a specific node by their affinity to it, and therefore ranking metrics such as NDCG or MRR are used. Moreover, TLP on weighted dynamic graphs operates in a discrete-time setting, while node affinity prediction operates in a continuous-time setting. Consequently, solutions for the former (Qin et al., 2023; Yang et al., 2019; Lei et al., 2019) operate on the full graph at each change, which may lead to unreasonable runtime if applied in the latter setting. In addition, upon each query, TLP on weighted dynamic graphs requires computing the full weighted adjacency matrix, while for node affinity prediction only the affinity scores between the queried node and the other nodes need to be computed. These factors are likely to hinder methods for TLP on weighted dynamic graphs from transferring well to node affinity prediction, and vice versa.
@@ 6 Conclusion
Limitations and future work
Although NAViS has shown strong performance, both theoretically, by generalizing common heuristics that excel in node affinity prediction, and empirically on TGB benchmarks, NAViS remains elementary, and further research in this direction is required. For example, NAViS currently utilizes a basic virtual global state based on recency selection from a global buffer to capture trends in dynamic graphs. Advanced aggregation schemes over the global buffer (e.g., attention-based mechanisms) and more sophisticated buffer-eviction strategies (e.g., non-deterministic eviction) may improve the modeling of global state and further enhance TGNN performance on node affinity prediction. In addition, NAViS computes node affinities as a linear combination of the input and current state, with coefficients in $[0,1]$ that are adaptive to both. Consequently, NAViS cannot, for example, represent certain non-linear functions. This limitation could be addressed by adding a third non-linear component with an associated coefficient such that all three coefficients sum to 1. We leave to future work the investigation of how to best combine linear and non-linear components to enable TGNNs to perform optimally on the node affinity task.
Acknowledgments
ME acknowledges support from the Israeli Ministry of Innovation, Science & Technology.
@@ Appendix A Datasets statistics and description
  • tgbn-reddit (Nadiri and Takes, 2022) models Reddit as a temporal bipartite graph linking users and subreddits. Nodes represent both entities, and a timestamped edge records a user’s post within a subreddit. The dataset covers activity from 2005 through 2019. The predictive objective is, for each user, to produce a next-week ranking of subreddits by expected engagement intensity.
@@ Appendix D Implementation Details
We initialized each linear layer in NAViS with values drawn uniformly from the range $\sqrt{d},\sqrt{d}]$, where $d$ is the input dimension of the layer. No normalization of the output vector $s$ were required as the ground true affinity scores in TGB are already normalized. The final loss we used is $\ell_{\mathrm{lambda}}+\mathcal{C}\ell_{\mathrm{Reg}}$, where $\mathcal{C}$ is a regularization coefficient. In the experiments we set $\mathcal{C}=1$. We adopted standard hyperparameters, consistent with prior work: batch size 200, Adam optimizer, and learning rate $10^{-4}$. Following the TGB protocol, all models were trained for 50 epochs, and the checkpoint with the best performance on the validation set was selected. We set the regularization margin to $\Delta=10^{-3}$. During training, NAViS computes the loss only over the top-20 affinities, ensuring that loss computation requires constant time and memory. We used a single NVIDIA GeForce RTX 3090 GPU and a single AMD Ryzen 9 7900X 12-Core Processor CPU.
Appendix E Additional results
Additional Evaluation Metrics
We additionally compare NAViS to the best performing baseline in each setting and report the results in terms of MRR (the average value of $\frac{1}{\mathrm{RANK}^{1}}$, where $\mathrm{RANK}^{1}$ is the predicted rank of the ground-truth top-scored node) and Recall@10 (the average number of top-10 ground-truth nodes ranked within the top-10 predicted scores). We report the results in Tables˜7, 8, 9 and 10. From the results, we observe that the performance gap between the baselines increases under the MRR metric but decreases under the Recall@10 metric. This can be explained by our objective, which penalizes mismatches on top-ranked ground-truth nodes (e.g., ranks 1–3) more strongly than on lower-ranked ones (e.g., ranks 8–10).
Table 7: MRR on TGB datasets (↑ higher is better). NAVIS is benchmarked against TGNNs that use all available graph messages. Boldface marks the best method.
Method tgbn-trade tgbn-genre tgbn-reddit tgbn-token
Val. Test Val. Test Val. Test Val. Test
[…]
Table 8: Recall on TGB datasets (↑ higher is better). NAVIS is benchmarked against TGNNs that use all available graph messages. Boldface marks the best method.
Method tgbn-trade tgbn-genre tgbn-reddit tgbn-token
Val. Test Val. Test Val. Test Val. Test
[…]
Table 9: MRR on TGB datasets using only previous ground-truth labels (↑ higher is better). This setting is suited for heuristics. Boldface marks the best method. Baselines have no standard deviation because they are pre-defined and deterministic.
Method tgbn-trade tgbn-genre tgbn-reddit tgbn-token
Val. Test Val. Test Val. Test Val. Test
Moving Avg 0.77 0.78 0.31 0.32 0.41 0.40 0.44 0.47
[…]
Table 10: Recall@10 on TGB datasets using only previous ground-truth labels (↑ higher is better). This setting is suited for heuristics. Boldface marks the best method. Baselines have no standard deviation because they are pre-defined and deterministic.
Method tgbn-trade tgbn-genre tgbn-reddit tgbn-token
Val. Test Val. Test Val. Test Val. Test
Moving Avg 0.74 0.73 0.29 0.30 0.25 0.24 0.30 0.29
[…]
Noisy data experiment
To examine the robustness of NAViS, we performed an additional experiment on noisy data. We used the tgbn-genre benchmark and, for each node that received its $N$-th update, we added an extra random update from that node to a random node and weighted the interaction with largest magnitude of affinity. We ran this experiment twice, once with $N=10$ and a second time with $N=20$, and did not observe a significant change in the performance of NAViS on the dataset. This experiment provides empirical evidence that NAViS is robust to noise such as short-term spikes.
What NAViS learns?
Table˜2, the Persistent Forecast baseline performs better than Moving Average on tgbn-trade, while on tgbn-genre the opposite holds. This means that tgbn-genre requires more memory to accurately perform node affinity prediction while in tgbn-trade one needs to rely more on the newly arrived data. To examine what NAVIS learns in these two regimes, we measured the average values of $z_{h}$ and $z_{s}$ from Equation˜14, which control the relative influence of the memory component. We expect lower values on tgbn-trade to resemble persistent forecast and larger values on tgbn-genre to resemble moving average. The measured values are $z_{h}=0.48,z_{s}=0.49$ for tgbn-trade and $z_{h}=0.90,z_{s}=0.84$ for tgbn-genre. This alignment between the learned $z_{h},z_{s}$ and the heuristic that performs better on each dataset provides empirical evidence that NAVIS learns the intended heuristic behavior in practice.
Additional SSM comparison
We further compared NAViS to an additional general SSM baseline, the S4 block (Gu et al., 2022), and report the results in Table˜11. The comparison was performed under the same ground-truth label setting as in our main experiments. Both our method and the S4 block utilized the same learning rate, batch size, and number of training epochs. The results in Table˜11 show that although the S4 block achieves relatively strong performance on the tested benchmark, NAViS still outperforms it, indicating that there remains substantial room for improving general SSM architectures on this benchmark.
Table 11: NDCG@10 on TGB datasets using only previous ground-truth labels (↑ higher is better). Boldface marks the best method.
Method tgbn-trade tgbn-genre
Val. Test Val. Test
[…]
Appendix F Ablation of Design Choices
@@
We further ablate on the regularization term of our loss, and performed an additional experiment on the TGB datasets where the regularization term is not included in the loss computation. We report the results in Table˜13.
@@
We also performed an empirical analysis on the effect of the hyperparameters of delta and batch size on the perfance of NAViS, and report the results inTable˜14. From Table˜14 we can see that on the tested datasets and batch size values has no effect on the performance of NAViS. In addion, choosing too small or too large values for delta can slightly hurt the performance of NAViS.
Table 14: Analysis of different batch sized and deltas, results are measured in NDCG@10 on test sets of datasets from TGB, averaged for 3 runs.
Hyperparameter tgbn-trade tgbn-genre
[…]
Additionally, we performed an analysis to examine the effect of the global buffer size and its aggregation scheme on the performance of NAViS, compared to the recency selection aggregation scheme used in previous experiments. We report the results in Table˜15. The results in Table˜15 show that changing the aggregation scheme or the size of the buffer has little to no effect on the existing implementation of NAViS. The slight drop in performance when the buffer size is 8 and the aggregation scheme is mean may be explained by the reliance on relatively old values in the buffer. We leave the exploration of new and more sophisticated virtual global state mechanisms (e.g., attention-based aggregation) for future work.
Buffer size Aggregation scheme tgbn-trade tgbn-genre
[…]
Appendix G Memory and Runtime analysis
We conducted a runtime and memory analysis to examine the efficiency of NAViS. In Table˜16, we report the number of parameters of the TGNN baselines and NAViS. NAViS not only requires the fewest parameters among the compared methods, but also scales well to large graphs due to the sparsification pipeline detailed in Section˜3.2.
Table 16: Number of parameters of TGNN baselines and NAViS, compared to the total number of nodes in each TGB benchmark for node affinity prediction.
Method tgbn-trade tgbn-genre tgbn-reddit tgbn-token
#Nodes 255 1505 11766 61756
DyGMamba 255125 257963 259998 -
DyGFormer 1027877 1030715 1032750 -
TGN 207655 233713 252398 283401
TGNv2 6433023 6565377 6660282 6817769
NAVIS 1280 2570 3495 5010
In Table˜17, we report a runtime comparison for a single training epoch and inference between NAViS, heuristics, and TGNN baselines. NAViS is more efficient than the TGNN baselines and has a runtime comparable to the heuristics. Since all methods use the same batch size, their throughput is proportional to their runtime. Since the heuristics do not contain learnable parameters they only require a single pass over the training data to compute node states before entering the validation and test phases. Other methods require multiple iterations over the training data (epochs) like standard deep learning models.
Table 17: Training and inference runtimes of the TGNN baselines, heuristics and NAViS on the tgbn-genre dataset.
Method Training runtime (sec) Inference runtime (sec)
Persistent Forecast / Moving Avg 34 6
DyGMamba 2080 381
DyGFormer 1440 304
TGN 170 33
TGNv2 130 36
NAVIS 46 8
Appendix H Theoretical remarks
In Theorem˜2 we showed that standard memory cells (RNN, LSTM, and GRU) cannot express the simple Persistent Forecast heuristic. The proof assumes that the input to the cells may be unbounded, and since $\tanh$ and other nonlinear activation functions have bounded images, this prevents the memory cells from expressing this heuristic. A natural question is what happens when the input $x$ is bounded, e.g., in $[0,1]$ or [1,1][-1,1]. Recall that persistent forecast is defined by $h_{i}=x_{i}$, while an RNN cell is defined by $h_{i}=\tanh(W_{h}h_{i-1}+W_{x}x_{i}+b)$. Denote $f(x_{i})=W_{h}h_{i-1}+W_{x}x_{i}+b$. To obtain $h_{i}=x_{i}$, we need $\tanh(f(x_{i}))=x_{i}$, i.e., $\tanh=f^{-1}$ on the relevant range by composition of functions. However, $f^{-1}$ must be affine, as the inverse of an affine function, which contradicts the fact that $\tanh$ is not affine. The same argument applies to $\tilde{h}$ in the GRU cell from Equation˜13. This also holds for any non-affine activation on [1,1][-1,1], such as ReLU or leaky ReLU. On $[0,1]$, ReLU is equal to the identity, and hence applying ReLU to inputs normalized to this range allows these memory cells to learn the Persistent Forecast heuristic.
Tjandra et al. (2024) state that There exists a formulation of TGNv2 that can represent persistent forecasting and moving average of order k. To prove this claim Tjandra et al. (2024) utilize permutation matrices, block matrix and a dedicated generator vector for the memory module of TGNv2. However, in practice TGNv2 is officially implemented with a GRU cell as the memory module. We showed in Theorem˜2 that GRU cell cannot express the simple Persistent Forecast heuristic.
In Theorem˜3 we proved that the cross-entropy loss is suboptimal by showing that there exist two rankings such that one ranks the elements identically to the ground truth while the other does not, yet under cross-entropy the latter achieves a smaller loss, contrary to what is desired in the task of node affinity prediction. We then argued that there exist infinitely many such examples since the cross-entropy loss is a continuous function. Consider the cross-entropy loss evaluated on the ground-truth order vector and a correctly ordered score vector as a function of the first entry of the correctly ordered vector (all other entries are fixed). Denote this function by $f(x)$. Let $L_{\mathrm{inc}}$ be the cross-entropy loss of a fixed incorrectly ordered vector, and define $\Delta\coloneqq f(x)-L_{\mathrm{inc}}>0$. By continuity of $f$, there exists $\delta>0$ such that for any $x^{\prime}$ satisfying $|x^{\prime}-x|<\delta$ we have $|f(x^{\prime})-f(x)|<\Delta/2$, which implies $f(x^{\prime})>L_{\mathrm{inc}}$. Since there are infinitely many such $x^{\prime}$, we obtain infinitely many correctly ordered score vectors that incur a larger cross-entropy loss than the incorrectly ordered vector.
In Section˜3.3, we introduced our loss function and claimed that the guarantee in Theorem˜3 does not apply to it. We now formalize and prove this claim. Let $\mathbf{y}$} denote the ground-truth affinity vector, and let ${\mathbf{s}}_{1}$ and ${\mathbf{s}}_{2}$ be two predicted logit vectors such that: (i) the ordering induced by ${\mathbf{s}}_{1}$ exactly matches the ground-truth node affinity scores, and (ii) ${\mathbf{s}}_{2}$ differs from ${\mathbf{s}}_{1}$ only in the logit of either the most-affinitive node or the least-affinitive node, thereby inducing an ordering different from that of ${\mathbf{s}}_{1}$. Since all logits in ${\mathbf{s}}_{2}$ are identical to those in ${\mathbf{s}}_{1}$ except for the most- or least-affinitive node, for any pair $(i,j)$ where neither $i$ nor $j$ is the most- or least-affinitive node, we have $\delta^{1}_{i,j}={\mathbf{s}}_{1}(i)-{\mathbf{s}}_{1}(j)=\delta^{2}_{i,j}={\mathbf{s}}_{2}(i)-{\mathbf{s}}_{2}(j)$. For pairs $(i,j)$ in which $i$ is the most-affinitive node or $j$ is the least-affinitive node, we obtain $\delta^{2}_{i,j}<\delta^{1}_{i,j}$, because the logit of the most-affinitive node can only decrease in ${\mathbf{s}}_{2}$ relative to ${\mathbf{s}}_{1}$, and the logit of the least-affinitive node can only increase. In both Equations˜3 and 18, smaller differences between correctly ranked affinity scores yield smaller loss values.